find the slope of 2x+4y=9
Found 2 solutions by palanisamy, richwmiller:
Answer by palanisamy(496)

THE EQUATION OF THE GIVEN LINE IS
2x+4y = 9
Slope of this line is m = (-coefficient of x/coefficient of y)
= -2/4
= -1/2

Answer by richwmiller(17219)

Putting this in slope intercept form
2x+4y=9
4y=9-2x
y=9/4-2/4*x
slope is -2/4 or -1/2